OpenStudy (anonymous):

FAN MEDAL INSTAGRAM Write the equation of the line that passes through the points (1, 7) and (5, 15) using function notation.

OpenStudy (anonymous):

please leave

OpenStudy (anonymous):

@MathGuyThatWantsHelp can u help me

OpenStudy (anonymous):

y- y1 = m (x -x1)

OpenStudy (anonymous):

Here r the answers y = 2x + 5 y = 4x + 8 f(x) = 2x + 5 f(x) = 4x + 8

OpenStudy (anonymous):

please someone help me

OpenStudy (anonymous):

m = y2-y1/x2-x1 lets suppose x1, y1 = (1,7) and x2, y2 = (5,15) then m (the slope) = (15-7)/5-1 = 8/4 = 2

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

so which one would it be i didn't really understand this concept during school @BPDlkeme234

OpenStudy (anonymous):

i need help!!!!!!!!!!!!!!!!111

OpenStudy (anonymous):

@BPDlkeme234

OpenStudy (anonymous):

please help me dude

OpenStudy (anonymous):

so, x1, y1 = (1,7) and (x2, y2) = (5, 15) this is known as the point-slope form of the equation. therefore the eq. of the line is : y-(7) = 2(x-1) y-7 = 2x -2

OpenStudy (anonymous):

now, if we add 2 to both sides we get: y-7 (+2) = 2x -2 (+2) y-5 =2x y= 2x + 5 is that one of your answers?

OpenStudy (anonymous):

no the Answers are y = 2x + 5 y = 4x + 8 f(x) = 2x + 5 f(x) = 4x + 8

OpenStudy (anonymous):

eh, yes it is equal to the first answer y=2x + 5